OpenStudy (anonymous):

1 and 1/3 times 5 and 5/8.

OpenStudy (anonymous):

thats incorrect i think...

OpenStudy (anonymous):

1 and 1/3 times 5 and 5/8. 1 and 1/3 = 4/3 5 and 5/8 = 45/8 4 45 1 15 15 1 ---- * ------ = ----- * ------ = ----- = 7 ---- 3 8 1 2 2 2

OpenStudy (anonymous):

ooopps, really sorry ;)

OpenStudy (anonymous):

Step 1 : Convert all numbers to fractions. Step 2 : 1 plus 1/3 = 4/3 ; 5 plus 5/8 = 45/8. Step 3 : Multiply across the top of both fractions : ie, 4 x 45 = 180. Step 4 : Multiply across the bottom of both fractions : ie, 3 x 8 = 24. Step 5 : Divide 180 by 24 = 7 and one half.